# DPER: Diffusion Prior Driven Neural Representation for Limited Angle and Sparse View CT Reconstruction

### v1

@@ Abstract
Limited-angle and sparse-view computed tomography (LACT and SVCT) are crucial for expanding the scope of X-ray CT applications. However, they face challenges due to incomplete data acquisition, resulting in diverse artifacts in the reconstructed CT images. Emerging implicit neural representation (INR) techniques, such as NeRF, NeAT, and NeRP, have shown promise in under-determined CT imaging reconstruction tasks. However, the unsupervised nature of INR architecture imposes limited constraints on the solution space, particularly for the highly ill-posed reconstruction task posed by LACT and ultra-SVCT. In this study, we introduce the Diffusion Prior Driven Neural Representation (DPER), an advanced unsupervised framework designed to address the exceptionally ill-posed CT reconstruction inverse problems. DPER adopts the Half Quadratic Splitting (HQS) algorithm to decompose the inverse problem into data fidelity and distribution prior sub-problems. The two sub-problems are respectively addressed by INR reconstruction scheme and pre-trained score-based diffusion model. This combination initially preserves the implicit image local consistency prior from INR. Additionally, it effectively augments the feasibility of the solution space for the inverse problem through the generative diffusion model, resulting in increased stability and precision in the solutions. We conduct comprehensive experiments to evaluate the performance of DPER on LACT and ultra-SVCT reconstruction with two public datasets (AAPM and LIDC). The results show that our method outperforms the state-of-the-art reconstruction methods on in-domain datasets, while achieving significant performance improvements on out-of-domain datasets.
@@ 1 Introduction
To improve the image reconstruction performance, numerous model-based iterative reconstruction (MBIR) methods [2, 3, 4, 5] have been investigated for solving inverse problems. MBIR problems are often expressed as an optimization challenge using weighted least squares, composed of both a data consistency term and a regularization term. Two of the most widely recognized regularization techniques are the total variation (TV) [6] and anisotropic TV (ATV) [7], known for artifact removal and edge preservation respectively. Besides, $\ell_{0}$-norm regularization is proposed to overcome the weaknesses of TV that over smoothing image structures[8]. Xu et al. proposed $\ell_{0}$ method using dictionary learning and image gradient $\ell_{0}$-norm to enhance the sharpness of reconstruction results[9]. Despite improving CT reconstructions, MBIR methods still struggle with removing severe artifacts (i.e., the artifacts in LACT and ultra-SVCT) and suffer from defects like sensitive hyper-parameters and hand-designed priors.
@@
Benefiting from data-driven priors, supervised deep learning (DL) approaches [10] yield state-of-the-art (SOTA) performance for the ill-posed CT reconstruction. For the SVCT task, convolutional neural networks (CNNs) are mainstream architectures [11, 12, 13, 14, 15]. For example, Jin et al. [11] proposed FBPConvNet, the first CNN-based SVCT model, which trains a U-Net to transform low-quality images into high-quality images. Regarding the greater ill-posed nature of the LACT problem, combining supervised neural networks with traditional model-based algorithms is a more effective scheme [16, 17, 18, 19]. For instance, DCAR [16] performs SART, an iterative optimization CT reconstruction method, on output images of a well-trained U-Net to further improve data consistency with the SV projections. The combination of the model-based method and CNN thus significantly improves the reconstruction performance. However, these supervised DL methods for SVCT and LACT face two common challenges: (1) The performance of supervised learning methods highly depends on the data distribution of the training dataset (i.e., a large-scale training dataset that covers more types of variations generally provides better performance). (2) Generalization of supervised models is susceptible to varying acquisition protocols (e.g., different SNR levels, projection views, and X-ray beam types), limiting their robustness in clinical application.
@@
Implicit neural representation (INR) is a novel unsupervised DL framework that demonstrated a significant potential for the SVCT problem [20, 21, 22, 23, 24, 25, 26, 27]. Technically, INR trains a multi-layer perceptron (MLP) to represent the reconstructed CT image as a continuous function that maps the spatial coordinates to the image intensities. Another method with similar functionality is the deep image prior (DIP) [28], which has been applied in CT, PET, and MR image reconstruction [29, 30, 31]. DIP operates under the assumption that the neural network architecture itself provides a regularization effect, and the network parameters should be estimated by fitting the measurement data without separate training. Compared with DIP, INR utilizes spatial coordinate embedding to achieve a continuous representation of the reconstructed image intensities. This approach exhibits enhanced capabilities in capturing high-frequency image details and proves to be more practical for incorporating various imaging forward model. By incorporating the CT imaging forward model (e.g., Radon transform), the MLP is optimized to approximate the SV projection data. The inherent consistency within INR imparts robust implicit regularization to constrain the solution space for the under-determined inverse problem [32]. Nonetheless, this regularization might hold less significance in scenarios that involve more severe data gaps, such as ultra-SVCT (e.g., less than 30 projection views) and LACT reconstruction. Thus, existing INR-based methods hardly produce satisfactory performance in the related tasks.
@@
Recently, score-based diffusion models have emerged as state-of-the-art generative models while also providing excellent generative priors to solve inverse problems in an unsupervised manner. Score-based diffusion models define a forward process that gradually perturbs the data with noise that transforms the data distribution into Gaussian. Through the reverse process, the diffusion models learn the data distribution by matching the gradient of the log density. Given a partial and corrupted measurement, the pre-trained diffusion models can sample a prior from estimated posterior probabilities. The prior facilitates constraining the space towards a feasible solution for ill-posed inverse problems. For inverse problems in medical imaging, there have been several methods based on diffusion models for SVCT, LACT and compressed sensing MRI (CS-MRI) that achieved notable progress [33, 34, 35, 36, 37]. Additionally, it has been demonstrated that a strong prior is essential for INR to achieve satisfactory results in these challenging scenarios [27]. Motivated by these developments, we suppose the diffusion models can serve as an implicit image prior incorporating within the INR optimization process for solving highly ill-posed inverse problems.
@@
In this work, we propose Diffusion Prior driven nEural Representation, an effective unsupervised framework to solve highly ill-posed inverse problems in CT reconstruction, including ultra-SVCT and LACT, namely DPER. To combat the ill-posedness in CT reconstruction, we incorporate the score-based diffusion models as a regularizer in INR optimization to constrain the space of feasible solutions. Following plug-and-play reconstruction methods, we decouple the data fidelity and data prior terms in inverse problem solvers and formulate them as two subproblems that are optimized alternatively. Specifically, in the data-fidelity subproblem, we propose a physical model-driven INR-based method. This approach allows us to explicitly model the image acquisition process during the fitting of measurement data. As a result, it leverages the advantages of both the physical model and the continuity prior, leading to a more comprehensive and accurate representation of the reconstructed images compared to traditional MBIR methods [27, 20, 25]. Regarding data prior, the score-based diffusion model samples a prior image by estimating posterior probabilities based on the reconstruction result from INR. Through alternating optimizations, the DPER arrives at a feasible solution while ensuring exceptional data consistency. To evaluate the effectiveness of the proposed method, we conduct experiments on SVCT and LACT reconstruction on public AAPM and LIDC datasets. The main contributions can be summarized as follows:
@@
    We propose DPER, a novel unsupervised approach that effectively solves ultra-SVCT and LACT reconstruction challenging problems.
@@
    We propose integrating the diffusion model as a generative prior into the INR reconstruction framework, significantly improving its capacity to tackle the highly ill-posed inverse problems.
@@
    We propose an effective alternating optimization strategy to enable an effective interaction between the diffusion model and INR, significantly improving reconstruction performance.
We conduct comprehensive evaluation experiments. The proposed model outperforms the SOTA DiffusionMBIR [35] for the LACT task and performs similarly for the ultra-SVCT task. For instance, PSNR improves about +2 dB for the LACT task of 90° scanning range and +5 dB for the LACT task of 120° scanning range.
@@ 2.1 Problem Formulation
Under SV and LA CT acquisition settings, the dimension of the measurement $\mathbf{y}$ is often much lower than that of the image $\mathbf{x}$ (i.e., $m\ll n$), resulting in the CT reconstruction task being highly ill-posed without further assumptions. A general scheme is thus to solve an optimization problem by adding suitable regularization for $\mathbf{x}$, which can be expressed as follows:
$\hat{\mathbf{x}}=\underset{\mathbf{x}}{\mathrm{arg}\min}\frac{1}{2}\left\|\mathbf{y}-\mathbf{A}\mathbf{x}\right\|^{2}+\lambda\cdot\mathcal{R}\left(\mathbf{x}\right),$ (2)
@@ 2.2 Implicit Neural Representation
The INR architectures usually consist of an encoding module (e.g., Fourier encoding [39] and hash encoding [40]) and an MLP network. The encoding module transforms low-dimensional coordinates into high-dimensional embeddings, significantly enhancing the fitting capacity of the subsequent MLP network. Due to the learning bias of the MLP network to low-frequency signals [32] and integration of the CT forward model, the INR imaging framework has shown significant potential in the SVCT reconstruction problem [20, 21, 22, 23, 24, 25, 26, 27].
@@ 2.3 Score-Based Diffusion Model
The score-based diffusion model is a generative model that progressively transforms pure noise to the target data distribution with a reverse diffusion process guided by the score function. It synthesizes samples by gradually reversing the stochastic process, introducing controlled noise while considering data distribution gradients. The diffusion process $\{\mathbf{x}(t)\}^{T}_{t=0}$ is indexed by a continuous time variable $t\in[0,T]$, such that $\mathbf{x}(0)\sim p_{0}$ and $\mathbf{x}(T)\sim p_{T}$, where $p_{0}$ is the data distribution and $p_{T}$ is the prior distribution (e.g., the Gaussian distribution). The forward diffusion process can be denoted by the solution of the forward stochastic differential equation (SDE):
@@
where the posterior is divided into $p_{t}(\mathbf{x})$ and $p_{t}(\mathbf{y}|\mathbf{x})$. Consequently, the unconditional pre-trained diffusion models can be harnessed for conditional synthesis, which enables us to regard diffusion model as a prior regularization term for solving inverse problems.
@@ 3.1 Overview
LACT and ultra-SVCT present highly under-determined inverse image reconstruction challenges. The image local consistency prior induced by INR hardly produces dependable solution space in such conditions [27, 20, 25]. In this study, we suggest integrating the diffusion model as a generative prior into the INR image reconstruction framework. The data distribution prior introduced by pre-trained score-based diffusion models serves as image regularization during the INR optimization process. This combination initially preserves the implicit image local consistency prior from INR. Additionally, it augments the feasibility of the solution space for the inverse problem through the diffusion model, resulting in increased stability and precision in the solutions.
Revisiting the ill-posed inverse problem as outlined in Eq. (2), if we reinterpret this issue within the framework of a diffusion model schema, it equates to resolving the Maximum A Posteriori (MAP) estimation problem at every timestep $t$:
$\hat{\mathbf{x}_{t}}=\underset{\mathbf{x}}{\mathrm{arg}\min}\frac{1}{2\sigma^{2}}\left\|\mathbf{y}-\mathbf{A}\mathbf{x}_{t}\right\|^{2}+\lambda\cdot\mathcal{R}\left(\mathbf{x_{t}}\right),$ (9)
@@
To solve Eq. (9) more efficiently, a widely used approach is decoupling the data and prior terms using the split-variable method. We leverage the Half Quadratic Splitting (HQS) [41], it first introduces an auxiliary variable $\mathbf{z}$, and the Eq. (2) is equivalent to the constrained optimization problem as:
@@
$\hat{\mathbf{x}_{t}}=\underset{\mathbf{x}_{t}}{\mathrm{arg}\min}\frac{1}{2\sigma^{2}}\left\|\mathbf{y}-\mathbf{Ax}_{t}\right\|^{2}+\lambda\cdot\mathcal{R}\left(\mathbf{z}_{t}\right),\ \operatorname{s.t.}\ \mathbf{z}_{t}=\mathbf{x}_{t},$ (10)
and Eq. (10) can be solved by minimizing the problem as follows:
$\mathcal{L}_{\mu}\left(\mathbf{x},\mathbf{z}\right)=\frac{1}{2\sigma^{2}}\left\|\mathbf{y}-\mathbf{Ax}\right\|^{2}+\lambda\cdot\mathcal{R}\left(\mathbf{z}\right)+\frac{\mu}{2}\left\|\mathbf{z}-\mathbf{x}\right\|^{2},$ (11)
@@
where $\mu$ is a penalty parameter. The problem can be addressed by iteratively solving the following two subproblems while keeping the rest of the variables fixed:
$\mathbf{x}_{t}=\underset{\mathbf{x}_{t}}{\mathrm{arg}\min}\left\|\mathbf{y}-\mathbf{A}\mathbf{x}_{t}\right\|^{2}+\mu\sigma_{i}^{2}\left\|\mathbf{x}-\mathbf{z}_{t-1}\right\|^{2}$ (12a)
@@
$\mathbf{z}_{t}=\underset{\mathbf{z}_{t}}{\mathrm{arg}\min}\frac{1}{2(\sqrt{\lambda/\mu})^{2}}\left\|\mathbf{z}_{t}-\mathbf{x}_{t}\right\|^{2}+\mathcal{R}\left(\mathbf{z}_{t}\right),$ (12b)
where Eq. (12a) defines the data-fidelity subproblem, concentrating on the identification of a proximal point in relation to $\mathbf{z}_{t-1}$. Meanwhile, the prior subproblem, as explicated in Eq. (12b), embodies a Gaussian denoising challenge, viewed through the lens of Bayesian analysis.
@@
Firstly, we describe how to solve the prior subproblem (as Eq. (12b)) using the score-based diffusion model under the HQS framework as Eq. (12). This Gaussian denoising problem involves deriving a noise-free $\mathbf{z}_{t}$ from a noisy $\mathbf{x}_{t}$ (i.e., obtain the $\mathbf{x}_{0|t}$), where the noise level is quantified by $\sqrt{\lambda/\mu}$. In alignment with the approach adopted in [42], the noise schedule of our Variance Exploding Stochastic Differential Equation (VE-SDE) diffusion model is applied, setting $\sqrt{\lambda/\mu}=\sigma_{t}$. Furthermore, we introduce $\nabla_{\mathbf{x}}\mathcal{R}\left(\mathbf{x}\right)=-\nabla_{\mathbf{x}}\log p_{t}\left(\mathbf{x}\right)=-\mathbf{s}_{\bm{\theta}^{*}}(\mathbf{x}_{t},t)$, we can rewrite Eq. (12b) as:
@@
which means that Eq. (12b) has been transformed into a Tweedie denoising process in VE-SDE diffusion schema. Then we can solve the data-fidelity subproblem to get a refined $\hat{\mathbf{x}}_{0|t}$ corrected with measurement $\mathbf{y}$. However, due to the highly undercharacterized nature of the problem, it is difficult to produce the desired result in a single iteration. To take full advantage of the multi-step sampling of the diffusion model, we follow [42] to add noise to the refined $\hat{\mathbf{x}}_{0|t}$ with noise level corresponding to timestep $t$. The operation mentioned above can be succinctly formulated as follows:
@@
$\hat{\mathbf{x}}_{0|t}=\underset{\mathbf{x}}{\mathrm{arg}\min}\left\|\mathbf{y}-\mathbf{A}\mathbf{x}\right\|^{2}+\mu\sigma_{i}^{2}\left\|\mathbf{x}-\mathbf{x}_{0|t}\right\|^{2},$ (15b)

### v2

@@ Abstract
Limited-angle and sparse-view computed tomography (LACT and SVCT) are crucial for expanding the scope of X-ray CT applications. However, they face challenges due to incomplete data acquisition, resulting in diverse artifacts in the reconstructed CT images. Emerging implicit neural representation (INR) techniques, such as NeRF, NeAT, and NeRP, have shown promise in under-determined CT imaging reconstruction tasks. However, the unsupervised nature of INR architecture imposes limited constraints on the solution space, particularly for the highly ill-posed reconstruction task posed by LACT and ultra-SVCT. In this study, we introduce the Diffusion Prior Driven Neural Representation (DPER), an advanced unsupervised framework designed to address the exceptionally ill-posed CT reconstruction inverse problems. DPER adopts the Half Quadratic Splitting (HQS) algorithm to decompose the inverse problem into data fidelity and distribution prior sub-problems. The two sub-problems are respectively addressed by INR reconstruction scheme and pre-trained score-based diffusion model. This combination first injects the implicit image local consistency prior from INR. Additionally, it effectively augments the feasibility of the solution space for the inverse problem through the generative diffusion model, resulting in increased stability and precision in the solutions. We conduct comprehensive experiments to evaluate the performance of DPER on LACT and ultra-SVCT reconstruction with two public datasets (AAPM and LIDC), an in-house clinical COVID-19 dataset and a public raw projection dataset created by Mayo Clinic. The results show that our method outperforms the state-of-the-art reconstruction methods on in-domain datasets, while achieving significant performance improvements on out-of-domain (OOD) datasets.
@@ 1 Introduction
To improve the image reconstruction performance, numerous model-based iterative reconstruction (MBIR) methods [2, 3, 4, 5] have been investigated for solving inverse problems. MBIR problems are often expressed as an optimization challenge using weighted least squares, composed of both a data consistency term and a regularization term. Typical regularization terms include total variation (TV) [6], anisotropic TV (ATV) [7], dictionary learning [8, 9, 10], and low-rank [11]. While these methods have demonstrated effectiveness, several limitations impede their widespread application. For example, TV tends to over-smooth image structures, whereas dictionary learning suffers from redundancy and high computational memory requirements. To address these issues, researchers have proposed several novel solutions. For instance, $\ell_{0}$-norm regularization [12, 9, 13] has been introduced to mitigate the image structure smoothing inherent in TV methods. Additionally, the convolutional analysis operator learning (CAOL) approach [14, 15] has been suggested as a potential solution to the high memory cost associated with dictionary learning. Despite the improvement in CT reconstructions, MBIR methods still struggled with removing severe artifacts (i.e., the artifacts in LACT and ultra-SVCT) and suffer from defects like sensitive hyper-parameters and hand-crafted priors.
@@
Benefiting from data-driven priors, supervised deep learning (DL) approaches [16] yield state-of-the-art (SOTA) performance for the ill-posed CT reconstruction. For the SVCT task, convolutional neural networks (CNNs) are mainstream architectures [17, 18, 19, 20, 21]. For example, Jin et al. [17] proposed FBPConvNet, the first CNN-based SVCT model, which trains a U-Net to transform low-quality images into high-quality images. In addition to direct image-to-image translation methods, combining supervised neural networks with the MBIR method is a more effective scheme [22, 23, 24, 25, 26]. For instance, DCAR [22] performs SART, an iterative optimization CT reconstruction method, on output images of a well-trained U-Net to further improve data consistency with the SV projections. Adler and Öktem  [27] proposed an approach that unrolls the proximal primal-dual optimization algorithm for tomographic reconstruction, wherein they replace the proximal operators with CNNs. Chen et al. [28] and Zhang et al. [29] introduced CNNs to learn the gradient of regularization terms in the gradient descent algorithm. In addition to image domain post-processing methods, Wu et al. [30] introduced a dual-domain method that extracts deep features from both the image acquisition and reconstruction domains, thereby leveraging the strengths of compressed sensing and deep learning techniques. The combination of the model-based method and CNNs significantly improves the reconstruction performance. However, these supervised DL methods for SVCT and LACT face two common challenges: (1) The performance of supervised learning methods highly depends on the data distribution of the training dataset (i.e., a large-scale training dataset that covers more types of variations generally provides better performance). (2) Generalization of supervised models is susceptible to varying acquisition protocols (e.g., different SNR levels, projection views, and X-ray beam types), limiting their robustness in clinical applications.
@@
Implicit neural representation (INR) is a novel unsupervised DL scheme that demonstrated a significant potential for the SVCT problem [31, 32, 33, 34, 35, 36, 37, 38]. Technically, INR trains a multi-layer perceptron (MLP) to represent the reconstructed CT image as a continuous function that maps the spatial coordinates to the image intensities. Another method with similar functionality is the deep image prior (DIP) [39], which has been applied in CT, PET, and MR image reconstruction [40, 41, 42]. DIP operates under the assumption that the neural network architecture inherently provides a regularization effect, and the network parameters are estimated by directly fitting the measurement data without requiring separate training. Compared to DIP, INR introduces spatial coordinate embedding to achieve a continuous representation of the reconstructed image intensities. This approach enhances INR’s capability to capture high-frequency image details and has been proven more practical for incorporating various imaging forward models. For instance, in SVCT reconstruction, the MLP is optimized by incorporating the CT imaging forward model (e.g., Radon transform) to approximate the SV projection data. The neural representation ensures inherent local image consistency, providing robust implicit regularization to constrain the solution space for the under-determined inverse problem [43]. Nonetheless, this regularization may be less effective in scenarios with more severe data gaps, such as ultra-SVCT (e.g., fewer than 30 projection views) and LACT reconstruction. As a result, existing INR-based methods hardly produce satisfactory performance in the related tasks.
@@
Recently, score-based diffusion models have emerged as state-of-the-art generative models while also providing excellent generative priors to solve inverse problems in an unsupervised manner. Score-based diffusion models define a forward process that gradually perturbs the data with noise that transforms the data distribution into Gaussian. Through the reverse process, the diffusion models learn the data distribution by matching the gradient of the log density. Pre-trained diffusion models can sample a prior from estimated posterior probabilities when given partial or corrupted measurements. The prior effectively constrains the solution space for challenging inverse problems. In medical imaging, diffusion model-based approaches have shown significant advancements in SVCT, LACT, and compressed sensing MRI [44, 45, 46, 47, 48, 49, 50]. Moreover, it has been shown that a robust prior is crucial for achieving satisfactory results in challenging scenarios with INR [38]. Inspired by these advancements, we propose integrating diffusion models as an implicit image prior within the INR optimization process to tackle highly ill-posed inverse problems.
@@
In this work, we propose DPER: Diffusion Prior driven nEural Representation, an effective unsupervised framework for solving highly ill-posed inverse problems in CT reconstruction, including ultra-SVCT and LACT. To address the ill-posed nature of CT reconstruction, we integrate score-based diffusion models as a regularizer in the INR optimization to constrain the solution space effectively. Following plug-and-play (PnP) [51, 52] reconstruction methods, we decouple the data fidelity and data prior terms in inverse problem solvers, formulating them as two subproblems optimized alternatively. Specifically, in the data-fidelity subproblem, we propose a physical model-driven INR-based method. This approach explicitly models the image acquisition process during the fitting of measurement data. As a result, it leverages the advantages of both the physical model and the neural representation continuity prior, resulting in a more comprehensive and accurate representation of the reconstructed images compared to traditional MBIR methods [38, 31, 36]. Regarding data prior, the score-based diffusion model samples a prior image by estimating posterior probabilities based on the reconstruction result from INR. Through alternating optimizations, DPER ensures that the reconstruction results stay within the target manifold of clean CT images, while keeping exceptional data consistency. To evaluate the effectiveness of the proposed method, we conduct experiments on SVCT and LACT reconstruction on public AAPM and LIDC datasets, an in-house clinical COVID-19 dataset and a public raw projection dataset. The main contributions can be summarized as follows:
@@
    We propose DPER, a novel unsupervised method that fully combines the respective advantages of diffusion modeling and INR, which can effectively solve ultra-SVCT and LACT reconstruction challenging problems.
@@
    We provide a new perspective for solving the medical imaging inverse problem via diffusion models. Innovatively, DPER, integrates diffusion models and INR as operators within a plug-and-play (PnP) framework for image restoration. This alternating iterative optimization process strictly adheres to the PnP framework, theoretically ensuring convergence to the optimal reconstruction result.
@@
    We provide a new insight of how to add data-driven regularizers to the INR frameworks. Our framework successfully and efficiently integrates the state-of-the-art diffusion model as a data prior to INR, creating the potential for broader applications of the INR framework.
We conduct comprehensive evaluation experiments, demonstrating that the proposed model significantly outperforms the state-of-the-art DiffusionMBIR [46] and MCG [45] in both the LACT and ultra-SVCT tasks. Notably, our model achieves an impressive +2 dB improvement in PSNR for the LACT task with a 90° scanning range and an outstanding nearly +5 dB improvement for the LACT task with the same setting on OOD data and real projection data.
@@ 2.1 Problem Formulation
Under SV and LA CT acquisition settings, the dimension of the measurement $\mathbf{y}$ is often much lower than that of the image $\mathbf{x}$ (i.e., $m\ll n$), resulting in the CT reconstruction task being highly ill-posed without further assumptions. This problem is often formulated as a maximum-a-posteriori (MAP) estimation [53], with the objective of maximizing the posterior probability:
$\hat{\mathbf{x}}$ $=\underset{\mathbf{x}}{\mathrm{arg}\max}\ p\left(\mathbf{x}\mid\mathbf{y}\right)$ (2)
$=\underset{\mathbf{x}}{\mathrm{arg}\min}-\log p\left(\mathbf{y}\mid\mathbf{x}\right)-\log p\left(\mathbf{x}\right),$ (3)
where the conditional probability $p\left(\mathbf{y}\mid\mathbf{x}\right)$ models the CT forward imaging process, and the prior distribution $p\left(\mathbf{x}\right)$ describes the probability distribution of the prior manifold. The solution to the MAP problem in Eq. (3) is generally formulated as a regularized inversion [54], which can be expressed as follows:
$\hat{\mathbf{x}}=\underset{\mathbf{x}}{\mathrm{arg}\min}\left\|\mathbf{y}-\mathbf{A}\mathbf{x}\right\|^{2}+\lambda\cdot\mathcal{R}\left(\mathbf{x}\right),$ (4)
@@ 2.2 Implicit Neural Representation
The INR architectures typically consist of an encoding module (e.g., Fourier encoding [56] and hash encoding [57]) and an MLP network. The encoding module transforms low-dimensional coordinates into high-dimensional embeddings, significantly enhancing the MLP network’s ability to fit high-frequency image content. Due to the MLP network’s inherent learning bias towards low-frequency signals [43] and the integration of the CT forward model, the INR imaging framework has demonstrated significant potential in addressing the SVCT reconstruction problem [31, 32, 33, 34, 35, 36, 37, 38].
@@ 2.3 Score-Based Diffusion Model
2.3.1 Background
The score-based diffusion model is a generative model that progressively transforms pure noise into the target data distribution through a reverse diffusion process guided by the score function. It synthesizes samples by gradually reversing the stochastic process, introducing controlled noise while considering data distribution gradients. The diffusion process $\{\mathbf{x}(t)\}^{T}_{t=0}$ is indexed by a continuous time variable $t\in[0,T]$, such that $\mathbf{x}(0)\sim p_{0}$ and $\mathbf{x}(T)\sim p_{T}$, where $p_{0}$ is the data distribution and $p_{T}$ is the prior distribution (e.g., the Gaussian distribution). The forward diffusion process can be denoted by the solution of the forward stochastic differential equation (SDE):
@@ 2.3.2 Diffusion model for inverse problems
where the posterior is divided into $p_{t}(\mathbf{x})$ and $p_{t}(\mathbf{y}|\mathbf{x})$. Consequently, the unconditional pre-trained diffusion models can be harnessed for conditional synthesis, which enables us to regard diffusion model as a prior regularization term for solving inverse problems. However, the noisy likelihood $\log p_{t}(\mathbf{y}|\mathbf{x})$ is generally intractable. The crux of the problem, therefore, lies in finding methods to approximate it or bypass direct calculation.
@@ 3.1 Overview
LACT and ultra-SVCT present highly under-determined inverse image reconstruction challenges. The image local consistency prior induced by INR hardly produces dependable solution space in such conditions [38, 31, 36]. In this study, we suggest integrating the diffusion model as a generative prior into the INR image reconstruction framework. The data distribution prior introduced by pre-trained score-based diffusion models serves as image regularization during the INR optimization process. This combination first preserves the implicit image local consistency prior from INR. Furthermore, it enhances the feasibility of the solution space for the inverse problem through the diffusion model, resulting in increased stability and precision in the solutions.
@@
Revisiting the ill-posed inverse problem as outlined in Eq. (4), which can be intuitively understood as finding a balance between the measurement manifold and the prior manifold. A widely used approach to solve it more efficiently is to decouple the data and prior terms using the split-variable method. Specifically, we leverage the Half Quadratic Splitting (HQS) [58] due to its straightforwardness and ease of implementation. The HQS algorithm first introduces an auxiliary variable $\mathbf{z}$, and the Eq. (4) is equivalent to the constrained optimization problem as:
@@
$\left\{\hat{\mathbf{x}},\hat{\mathbf{z}}\right\}=\underset{\mathbf{x},\mathbf{z}}{\mathrm{arg}\min}\left\|\mathbf{y}-\mathbf{Ax}\right\|^{2}+\lambda\cdot\mathcal{R}\left(\mathbf{z}\right),\ \operatorname{s.t.}\ \mathbf{z}=\mathbf{x},$ (11)
where the constraint is given by $\mathbf{z}=\mathbf{x}$. The HQS-based penalty method seeks to minimize the following cost function:
$\mathcal{L}_{\mu}\left(\mathbf{x},\mathbf{z}\right)=\left\|\mathbf{y}-\mathbf{Ax}\right\|^{2}+\lambda\cdot\mathcal{R}\left(\mathbf{z}\right)+\frac{\mu}{2}\left\|\mathbf{z}-\mathbf{x}\right\|^{2},$ (12)
@@
where $\frac{\mu}{2}\left\|\mathbf{z}-\mathbf{x}\right\|^{2}$ is a penalty term that penalizes the deviation of $\mathbf{z}$ from $\mathbf{x}$. The parameter $\mu$ controls the strength of this penalty. As $\mu$ increases, the penalty term becomes more dominant, leading $\mathbf{x}$ to approach $\mathbf{z}$ more precisely. The problem can be addressed by iteratively solving the following two subproblems while keeping the rest of the variables fixed:
$\mathbf{x}_{t}=\underset{\mathbf{x}}{\mathrm{arg}\min}\left\|\mathbf{y}-\mathbf{A}\mathbf{x}\right\|^{2}+\frac{\mu}{2}\left\|\mathbf{x}-\mathbf{z}_{t-1}\right\|^{2}$ (13a)
@@
$\mathbf{z}_{t}=\underset{\mathbf{z}}{\mathrm{arg}\min}\frac{1}{2(\sqrt{\lambda/\mu})^{2}}\left\|\mathbf{z}-\mathbf{x}_{t}\right\|^{2}+\mathcal{R}\left(\mathbf{z}\right),$ (13b)
since then, the original problem has been transformed into the standard plug-and-play image restoration (IR) framework [52]. Eq. (13a) defines the data-fidelity subproblem, concentrating on the identification of a proximal point in relation to $\mathbf{z}_{t-1}$. Meanwhile, the prior subproblem in Eq. (13b) can be interpreted as a denoising problem with additive white Gaussian noise (AWGN) of variance $(\sqrt{\lambda/\mu})^{2}$ [51]. To ensure that $\mathbf{x}_{t}$ and $\mathbf{z}_{t}$ in the iterative framework converge to a fixed point (i.e., the optimal solution to the original problem), $\mu$ needs to be gradually increased with iterations towards infinity ($\mu\rightarrow\infty$)[59]. For a detailed discussion on the theoretical convergence of plug-and-play IR, please refer to [54]
Firstly, we describe how to solve the prior subproblem (as Eq. (13b)) using the score-based diffusion model under the HQS framework in Eq. (13). Leverage the wisdom of PnP, the proximal operator in Eq. (13b) can be replaced by a general denoiser $\mathcal{D}\left(\cdot,\sigma^{2}\right):\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ as:
$\mathbf{z}_{t}=\mathcal{D}\left(\mathbf{x}_{t},(\sqrt{\lambda/\mu})^{2}\right),$ (14)
this AWGN denoising problem involves deriving a noise-free $\mathbf{z}_{t}$ from a noisy $\mathbf{x}_{t}$ (i.e., obtain the $\mathbf{x}_{0|t}$), where the noise level is quantified by $\sqrt{\lambda/\mu}$. By introducing the Tweedie’s formula [60], we can rewrite Eq. (13b) as:
$\mathbf{z}_{t}=\mathbf{x}_{0|t}$ $=\mathcal{D}\left(\mathbf{x}_{t},(\sqrt{\lambda/\mu})^{2}\right)$ (15)
$=\mathbb{E}\left[\mathbf{z}_{t}|\mathbf{x}_{t}\right]$ (16)
$=\mathbf{x}_{t}+\left(\sqrt{\lambda/\mu}\right)^{2}\nabla\log p\left(\mathbf{x}_{t}\right)$ (17)
$\approx\mathbf{x}_{t}+\left(\sqrt{\lambda/\mu}\right)^{2}\mathbf{s}_{\theta}\left(\mathbf{x}_{t},t\right).$ (18)
@@
In alignment with the approach adopted in [61], the noise schedule of our Variance Exploding Stochastic Differential Equation (VE-SDE) diffusion model is applied, setting $(\sqrt{\lambda/\mu})^{2}=\sigma_{t}^{2}$. Furthermore, we can rewrite Eq. (13b) as:
@@
where $\sigma_{t}^{2}$ is the noise level corresponding to timestep $t$ of the diffusion model. This means that Eq. (13b) has been transformed into a Tweedie denoising process in VE-SDE diffusion schema. We then proceed from $\mathbf{z}_{t}$ (i.e., $\mathbf{x}_{0|t}$), to solve the data-fidelity subproblem shown in Eq. (13a) to a refined $\hat{\mathbf{x}}_{0|t}$ corrected with measurement $\mathbf{y}$. Note that we have $\mu=\lambda/\sigma_{t}^{2}$, then the data-fidelity subproblem can be rewritten to:
$\hat{\mathbf{x}}_{0|t}=\underset{\mathbf{x}}{\mathrm{arg}\min}\left\|\mathbf{y}-\mathbf{A}\mathbf{x}\right\|^{2}+\frac{\lambda}{2\sigma_{t}^{2}}\left\|\mathbf{x}-\mathbf{x}_{0|t}\right\|^{2},$ (20)
where $\mu$ is a scaling factor hyper-parameter.
However, due to the highly undercharacterized nature of the problem, it is difficult to produce the desired result in a single iteration. To take full advantage of the multi-step sampling of the diffusion model, we follow [61] to add noise to the refined $\hat{\mathbf{x}}_{0|t}$ with noise level corresponding to timestep $t$. The operation mentioned above can be succinctly formulated as follows:
@@
$\hat{\mathbf{x}}_{0|t}=\underset{\mathbf{x}}{\mathrm{arg}\min}\left\|\mathbf{y}-\mathbf{A}\mathbf{x}\right\|^{2}+\frac{\lambda}{2\sigma_{t}^{2}}\left\|\mathbf{x}-\mathbf{x}_{0|t}\right\|^{2},$ (22b)